OpenStudy (anonymous):

5(x + 10) -25 = 5x +25 I need help on finding how many solutions. So how would I go about doing so with the equation above?

OpenStudy (anonymous):

5x + 50 - 25 = 5x + 25 then 5x-5x = 0 So your answer is 0

hero (hero):

5x+50 -25 = 5x + 25 5x + 25 = 5x + 25 Infinite Solutions

OpenStudy (anonymous):

^Agreed, forgot that.

hero (hero):

You're supposed to stop when both sides are equal, not keep going